OpenStudy (anonymous):

Prove: If n^3 ≡n (mod 6)

OpenStudy (inkyvoyd):

I remember this. this was annoying

ganeshie8 (ganeshie8):

Hint : \[n^3-n = n(n^2-1)=(n-1)n(n+1)\]

ganeshie8 (ganeshie8):

next, recall that any product of \(n\) consecutive integers is divisible by \(n!\) not that annoying :)